OpenStudy (toxicsugar22):

Molarity of standardized NaOH solution is 0.1012 final buret reading, mL IS 31.21 INITIAL buret reading, mL IS 15.72 VOLUME OF TITRANT, mL is 15.49 Moles of NaOH used is 1.568*10^-3 Find Moles of HC2H302 neutralized by the NaOH?

OpenStudy (abb0t):

Use: \(\sf \color{blue}{M_1V_1=M_2V_2}\)

OpenStudy (abb0t):

@nincompoop

OpenStudy (natsent06):

?

OpenStudy (anonymous):

@LunyMoony what'cha need?

OpenStudy (anonymous):

Hmmm I can try (=

OpenStudy (anonymous):

ok give me a minute (=

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Maybe This Will Help You: moles of NaOH = M*V = 0.1 * 0.00413 = 0.000413 mol moles of acetic acid = moles of NaOH = 0.000413 mol grams of acetic acid = moles * MM = 0.000413 * 60 = 0.0248 g %(m/v) = (g / mL) *100 = (0.0248 / 5) *100 = 4.96 %

OpenStudy (anonymous):

OMG THIS IS SO HARD! )=

OpenStudy (toxicsugar22):

I think we Use: M1V1=M2V2 equation

OpenStudy (anonymous):

@AuroraB

OpenStudy (anonymous):

@math

OpenStudy (anonymous):

@moderator

OpenStudy (aurorab):

HAHA yeahhhhhhhhh i can't do this... sorry... /:

OpenStudy (anonymous):

I'm not there yet sorry

OpenStudy (anonymous):

I know )=

OpenStudy (toxicsugar22):

Molarity of standardized NaOH solution is 0.1012 final buret reading, mL IS 31.21 INITIAL buret reading, mL IS 15.72 VOLUME OF TITRANT, mL is 15.49 Moles of NaOH used is 1.568*10^-3 Find Moles of HC2H302 neutralized by the NaOH?

OpenStudy (toxicsugar22):

the volume of diluted vinegar used in each titration is 20 ml .02 L

OpenStudy (cuanchi):

do you need to find the molarity of the acetic acid? you have to divide the number of moles by the volume in liters. 1.568*10^-3 /0.2= 7.84 x 10^-4 M

OpenStudy (toxicsugar22):

NO I HAVE TO FIND moles of HC2H302 NEUTRALIZED BY THE NaOH

OpenStudy (toxicsugar22):

IT THE ANSWER 1.568*10^-3

OpenStudy (cuanchi):

yes that the answer and already told you how to do the calculations it is that clear now?

OpenStudy (toxicsugar22):

YES

OpenStudy (toxicsugar22):

NOW I have already ruined my lab report. how do i print out a new one that is blank

OpenStudy (toxicsugar22):

like is there a trick

OpenStudy (toxicsugar22):

i want a new sheet of my lab report as i messed up the originla one

OpenStudy (toxicsugar22):

@abb0t is he right